OpenStudy (anonymous):

What is the sum of the arithmetic sequence 22, 13, 4 … if there are 28 terms?

OpenStudy (cwrw238):

sum n terms = (n/2)[2a + (n-1)d] n = no. terms, a = ist term and d = common difference

OpenStudy (anonymous):

n=28 ,a=22,d=-9, ________________________ now subs. in the following rule sum n terms = (n/2)[2a + (n-1)d]

OpenStudy (anonymous):

thnksss!

OpenStudy (anonymous):

your welcome :) ,any time ..